OpenStudy (anonymous):

Can someone help me solve this problem: h(t) = (square root of t) times (1-t^2)

OpenStudy (anonymous):

someone's actually solved this problem for me fbeofre, there is stil a certain part of this where i'm oconfused

OpenStudy (anonymous):

what exactly do you need done with this problem?

OpenStudy (anonymous):

Oh sorry, I need to take the derivative of this function. I've gotten to this part, I just don't know how to get the answer from here: \f'(x) =[(-2t)(\sqrt{t})+1-t ^{2}/2\sqrt{t}\]

OpenStudy (anonymous):

f'(x)=\[(-2t)(\sqrt{t})+1-t ^{2}/2\sqrt{t}\]

OpenStudy (anonymous):

h(t)=\[\sqrt{t}\](1+t)(1-t)

OpenStudy (anonymous):

here i will give you a website that is very helpful ans hows the steps. IF you still need help ask.

OpenStudy (anonymous):

yeah someone actually showed me that site the last time I used this service. The problem is actually this: http://www.wolframalpha.com/input/?i=%28t%29%5E1%2F2%281-t%5E2%29. I think I've almost figured out how to get the answer, just need to figure fix my mistakes.

OpenStudy (anonymous):

click show steps

OpenStudy (anonymous):

I'm stuck on the very last part. I know from there you would then have to get the lcd for each side, which would be \[2\sqrt{t}\]. and then times that on the left side so it would be \[2\sqrt{t} \times (-2t \sqrt{t})+1-t ^{2}/2\sqrt{t}\]

OpenStudy (anonymous):

What I'm confused about is, should I change the square root of t to t^1/2 and then take 2 square root of t times -2 square root of t and add the two exponents, 1/2 or just multiply it with the two square roots?